OpenStudy (anonymous):

carpet for a 12 ft by 8 ft room costs $300. find cost to carpet a 48 ft by 32 ft room.

OpenStudy (anonymous):

1200$ :)

OpenStudy (anonymous):

48x32/12x8 = x/$300 x=4*300=$1200

OpenStudy (anonymous):

thanks but these answers are wrong :)

OpenStudy (anonymous):

find area 12*8=96 ft^2 costs $300 1ft^2 costs $300/96 therefore 48*32=1536 ft^2 costs $300*1536/96=$4800

OpenStudy (turingtest):

Just to confirm, password1 gave the correct answer.